OpenStudy (wwe123):

if the equation x^2+px+q=0 and x^2+rs+s=0 have a common roots,prove that it is equal to

OpenStudy (wwe123):

\[\frac{ ps-rq }{ q-s } \] as well as \[\frac{ q-s }{ r-p }\] Find the remaining roots of the equations

OpenStudy (anonymous):

I think it is x^2+rx+s=0

jhonyy9 (jhonyy9):

the second equation is right,correct sure or you have missed anything ?

OpenStudy (wwe123):

yes @sauravshakya

jhonyy9 (jhonyy9):

how do you know if x1 and x2 are roots of a quadratic equation ax^2 +bx +c=0 how can you writing ,what will be x1 *x2 = and x1 + x2 = can you answer this using a,b and c ?

OpenStudy (anonymous):

Well roots of x^2+px+q=0 are 

OpenStudy (anonymous):

Can u tell me what are roots of x^2+rx+s=0

OpenStudy (wwe123):